the radius of yhe planet saturn is 5.87×10^7m and it's mass is 5.86×10^26kg. find the density of its material by g/m find it's surface area by m2 (4πr^2)
1
Expert's answer
2019-08-26T10:43:19-0400

The density by definition


"\rho=\frac{m}{V}=\frac{m}{4/3\pi R^3}"

"=\frac{5.86\times 10^{26}\:\rm{kg}}{4/3\pi (5.87\times 10^7\:\rm{m})^3}=692\:\rm{kg/m^3}=0.692\:\rm{g/cm^3}"

The surface area of the planet


"A=4\pi R^2=4\pi\times (5.87\times 10^7)^2=4.33\times 10^{16}\:\rm{m^2}"